OpenStudy (anonymous):

Find the following quotient. (Divide.) 28/ -4

OpenStudy (luigi0210):

Uhm, how many times does 4 go into 28?

OpenStudy (luigi0210):

28/-4 -(28/4) -(7/1) -7 ?

OpenStudy (anonymous):

I TRIED working it too.. i made a mistake and got 7

OpenStudy (anonymous):

and its correct that ist 7

OpenStudy (mathstudent55):

When you multiply or divide two numbers with the same sign (both positive or both negative) the answer is positive. Otherwise, the answer is negative. Divide the numbers ignoring the signs, then use the above rule to determine which sign to give the answer. For example: 32/8 = 4 -32/-8 = 4 32/-8 = -4 -32/8 = -4